# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

### v1

@@ 8.2. Proof of Theorem 3
holds for $t=0,\cdots,k$, where $R$ is the constant in Lemma 7 and $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (36) when $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CdR$ in (38) when $\sigma$ satisfies Asssumption 5.
Thanks to Corollary 3, it is sufficient to prove that Condition 3 also holds for $t=k+1$. For readability, we defer the proof of Corollary 3 to the end of this section. In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary 3 leads to Theorem 3.
Proof of Theorem 3.
We let $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (36) when $\sigma$ is the $\text{ReLU}^{3}$ activation function and let $R^{{}^{\prime\prime}}=CdR$ in (37) when $\sigma$ satisfies Assumption 5.
First, set $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$, then from Lemma 7 we have $\|\bm{J}(t)-\bm{J}(0)\|_{2}\leq\frac{\sqrt{\lambda_{0}}}{3}$, thus $\lambda_{min}(\bm{J}(t))\geq\frac{2\sqrt{\lambda_{0}}}{3}$ and $\lambda_{min}(\bm{H}(t))\geq\frac{4\lambda_{0}}{9}$ for $t=0,\cdots,k$.
From the updating role of NGD, we have
@@
$\left[\bm{J}(t)^{T}\right]_{r}=\left(\frac{\partial s_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial s_{n_{1}}(t)}{\partial\bm{w}_{r}},\frac{\partial h_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial h_{n_{2}}(t)}{\partial\bm{w}_{r}}\right).$
@@
$\leq\frac{9\eta}{4\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{2}\sqrt{L(t)}$
$\leq\frac{9\eta}{4\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{F}\sqrt{L(t)}$
$=\frac{9\eta}{4\lambda_{0}}\sqrt{\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}}\sqrt{L(t)}$
$\lesssim\frac{4\eta}{9\lambda_{0}}\sqrt{\frac{B^{4}+1}{m}}\sqrt{L(t)}$
@@
where the last inequality is due to Corollary 3.
@@
$\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}.$
Note that although $C$ is a universal constant, it still depends on the activation function.
Now, when $R^{{}^{\prime}}\leq 1$, we can deduce that Condition 3 also holds for $t=k+1$. Thus, it remains only to derive the requirement for $m$.
Recall that we need $m$ to satisfy that $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$.
(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CM\sqrt{R}\leq\frac{\sqrt{\lambda_{0}}}{3}$, wihch implies that $R\lesssim\frac{\lambda_{0}}{M^{2}}$. Then $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ imples that
$m=\Omega\left(\frac{M^{4}B^{4}L(0)}{\lambda_{0}^{4}}\right).$
From Lemma 11, we can deduce that
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
(2) When $\sigma$ satisfies Assumption 5, we have that
$R\lesssim\frac{\sqrt{\lambda_{0}}}{d},R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R.$
From Lemma 5 in [23], we know
$L(0)\lesssim d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right).$
Thus, we can deduce that
Proof of Corollary 3.
Similar as before, when $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$ , we have $\lambda_{min}(\bm{J}(t))\geq\frac{2\sqrt{\lambda_{0}}}{3}$ and then $\lambda_{min}(\bm{H}(t))\geq\frac{4}{9}\lambda_{0}$ for $t=0,\cdots,k$.
@@
where the second equality is from fundamental theorem of calculus and $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)=\bm{w}(t)-s\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(t)$.
Note that $\frac{\partial\bm{u}(\bm{w}(t))}{\partial\bm{w}}=\bm{J}(t)$, thus $\bm{I}_{1}(t)=\eta\bm{u}(t)$. Plugging it into (85) yields that
@@
$\lesssim\frac{\eta\sqrt{L(t)}}{\sqrt{\lambda_{0}}}R^{{}^{\prime\prime}}$
@@
$\leq\left[(1-\eta)^{2}+\frac{C^{2}\eta^{2}(R^{{}^{\prime\prime}})^{2}}{\lambda_{0}}+2(1-\eta)\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\right]\|u(t)\|_{2}^{2}$
@@
where the last inequality follows from the requirement
$2\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\leq\frac{\eta}{4},$
i.e., $R^{{}^{\prime\prime}}\lesssim\sqrt{\lambda_{0}}$.
Therefore, we can deduce that $\|\bm{u}(t)\|_{2}^{2}\leq(1-\eta)^{t}\|\bm{u}(0)\|_{2}^{2}$ holds for $t=0,\cdots,k$.
8.3. Proof of Corollary 1
@@
As $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)$, then for any $r\in[m]$, we have $\|\bm{w}_{r}(s)\|_{2}\leq s\|\bm{w}_{r}(t+1)\|_{2}+(1-s)\|\bm{w}_{r}(t)\|_{2}\leq B$.
When $\sigma$ is smooth, we can deduce that for any $r\in[m]$,
@@
Plugging it into (89), we have
@@
$\lesssim\frac{\sqrt{L(t)}}{\sqrt{\lambda_{0}}}\frac{\eta B^{4}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}$
@@
holds for $t\in\mathbb{N}$, where $C$ is a universal constant. ∎
@@ 9. Auxiliary Lemmas
Lemma 8 (Theorem 3.1 in [24]).
If $X_{1},\cdots,X_{n}$ are independent mean zero random variables with $\|X_{i}\|_{\psi_{\alpha}}<\infty$ for all $1\leq i\leq n$ and some $\alpha>0$, then for any vector $a=(a_{1},\cdots,a_{n})\in\mathbb{R}^{n}$, the following holds ture:
$P(|\sum\limits_{i=1}^{n}a_{i}X_{i}|\geq 2eC(\alpha)\|b\|_{2}\sqrt{t}+2eL_{n}^{*}(\alpha)t^{1/\alpha}\|b\|_{\beta(\alpha)})\leq 2e^{-t},\ for\ all\ t\geq 0,$
@@
and for $\beta(\alpha)=\infty$ when $\alpha\geq 1$ and $\beta(\alpha)=\alpha/(\alpha-1)$ when $\alpha>1$,
@@
Lemma 9 (Bernstein inequality, Theorem 3.1.7 in [25]).
@@
Lemma 10.
For $10<\delta<$ 1, with probability at least $1-\delta$, we have
$\|\bm{y}-\bm{u}(0)\|_{2}^{2}=\mathcal{O}\left(n\log\left(\frac{n}{\delta}\right)\right).$
Proof.
First, from Cauchy’s inequality, we have
$\|\bm{y}-\bm{u}(0)\|_{2}^{2}$ $=\sum\limits_{i=1}^{n}\left(y_{i}-\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}$
$\leq\sum\limits_{i=1}^{n}2y_{i}^{2}+2\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}$
$=2\sum\limits_{i=1}^{n}y_{i}^{2}+2\sum\limits_{i=1}^{n}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}.$
Since $\bm{w}_{r}(0)^{T}\bm{x}_{i}\sim\mathcal{N}(0,\|\bm{x}_{i}\|_{2})$ and $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$, we can deduce that
$\|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\|_{\psi_{2}}\leq\|\bm{w}_{r}(0)^{T}\bm{x}_{i}\|_{\psi_{2}}=\mathcal{O}(1)$
holds for all $r\in[m]$ and $i\in[n]$.
Thus for any fixed $i\in[n]$, applying Lemma 8 yields that with probability at least $1-2e^{-t}$,
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right|\leq C\sqrt{t},$
where $C$ is a universal constant. By taking union bound on all $i\in[n]$, the above inequality holds for all $i\in[n]$ with probability at least $1-2ne^{-t}$.
Let $2ne^{-t}=\delta$, i.e., $t=\log\left(\frac{2n}{\delta}\right)$, then we have
$\|\bm{y}-\bm{u}(0)\|_{2}^{2}\lesssim n+nt\lesssim n\log\left(\frac{n}{\delta}\right).$
Lemma 11.
@@
$L(0)=\left\|\begin{pmatrix}\bm{s}(0)\\ \bm{h}(0)\end{pmatrix}\right\|_{2}^{2}=\mathcal{O}\left(d^{6}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
@@
$\leq\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right)+\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}f^{2}(x_{p})$

### v3

@@ 10. Proof of Section 4
holds for $t=0,\cdots,k$, where $R$ is the constant in Lemma 4.5 and $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (16) when $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CdR$ in (18) when $\sigma$ satisfies Assumption 4.3.
Thanks to Corollary C.1, it is sufficient to prove that Condition 3 also holds for $t=k+1$. For readability, we defer the proof of Corollary C.1 to the end of this section. In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary C.1 leads to Theorem 4.7.
Sketch Proof of Theorem 4.7.
First, let $\bm{u}(t)=\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}$, then from the updating formula of NGD (11), we have
$\bm{u}(t+1)-\bm{u}(t)$ (79)
[…]
$:=\bm{I}_{1}(t)+\bm{I}_{2}(t),$
where the second equality is from the fundamental theorem of calculus and $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)=\bm{w}(t)-s\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(t)$.
In the proof, we assume that Condition 3 holds for $t=0,\cdots,k$. Then from Corollary C.1, to prove Theorem 4.7, it suffices to demonstrate that this condition also holds for $t=k+1$. Here, we primarily explain the process from Condition 3 to Corollary C.1, while other content is placed in the following full proof of Theorem 4.7.
Note that $\frac{\partial\bm{u}(\bm{w}(t))}{\partial\bm{w}}=\bm{J}(t)$, thus $\bm{I}_{1}(t)=\eta\bm{u}(t)$. Plugging this into (79) yields that
$\bm{u}(t+1)=(1-\eta)\bm{u}(t)+\bm{I}_{2}(t).$ (80)
From the above equation, we can see the difference between NGD and GD. Recall that the iteration formula for GD is
$\bm{u}(t+1)=(1-\eta\bm{H}(t))\bm{u}(t)+\bm{I}_{1}(t).$
Precisely because of this, the convergence rate of GD is inevitably influenced by $\lambda_{0}$ , whereas that of NGD is not.
From the stability of the Jacobian matrix, we can deduce that $\|\bm{I}_{2}(t)\|_{2}=\mathcal{O}(\eta\|\bm{u}(t)\|_{2}/\sqrt{m})$. Plugging this into (80) yields that
$\|\bm{u}(t+1)\|_{2}^{2}$ (81)
$\leq\|(1-\eta)\bm{u}(t)\|_{2}^{2}+\|\bm{I}_{2}(t)\|_{2}^{2}+2(1-\eta)\|\bm{u}(t)\|_{2}\|\bm{I}_{2}(t)\|_{2}$
$=\left((1-\eta)^{2}+\mathcal{O}\left(\frac{\eta^{2}}{m}\right)+2(1-\eta)\mathcal{O}\left(\frac{\eta}{\sqrt{m}}\right)\right)\|\bm{u}(t)\|_{2}^{2}$
$\leq(1-\eta)\|\bm{u}(t)\|_{2}^{2},$
where the last inequality holds if $m$ is large enough. ∎
Full Proof of Theorem 4.7.
Recall that we let $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (16) when $\sigma$ is the $\text{ReLU}^{3}$ activation function and let $R^{{}^{\prime\prime}}=CdR$ in (18) when $\sigma$ satisfies Assumption 4.3.
First, we can set $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{3\lambda_{0}}}{6}$, since $R^{{}^{\prime\prime}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$. Then from Lemma 4.5 we have $\|\bm{J}(t)-\bm{J}(0)\|_{2}\leq\frac{\sqrt{3\lambda_{0}}}{6}$, thus
$\sigma_{min}(\bm{J}(t))\geq\sigma_{min}(\bm{J}(0))-\|\bm{J}(t)-\bm{J}(0)\|_{2}\geq\frac{\sqrt{3\lambda_{0}}}{2}-\frac{\sqrt{3\lambda_{0}}}{6}=\frac{\sqrt{3\lambda_{0}}}{3}$
and then $\lambda_{min}(\bm{H}(t))\geq\frac{\lambda_{0}}{3}$ for $t=0,\cdots,k$, where $\sigma_{min}(\cdot)$ denotes the least singular value.
From the updating rule of NGD, we have
@@
$\left[\bm{J}(t)^{T}\right]_{r}=\left[\frac{\partial s_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial s_{n_{1}}(t)}{\partial\bm{w}_{r}},\frac{\partial h_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial h_{n_{2}}(t)}{\partial\bm{w}_{r}}\right].$
@@
$\leq\frac{3\eta}{\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{2}\sqrt{L(t)}$
$\leq\frac{3\eta}{\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{F}\sqrt{L(t)}$
$=\frac{3\eta}{\lambda_{0}}\sqrt{\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}}\sqrt{L(t)}$
$\lesssim\frac{\eta}{\lambda_{0}}\sqrt{\frac{B^{4}+1}{m}}\sqrt{L(t)}$
@@
where the last inequality is due to Corollary C.1.
@@
$\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}},$
where $C$ is a universal constant.
Now, when $R^{{}^{\prime}}\leq 1$, we can deduce that $\|\bm{w}_{r}(k+1)\|_{2}\leq B$, implying that Condition 3 also holds for $t=k+1$. Thus, it remains only to derive the requirement for $m$.
Recall that we need $m$ to satisfy that $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$.
(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, in Corollary C.1, $R^{{}^{\prime\prime}}=CM\sqrt{R}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$, implying that $R\lesssim\frac{(1-\eta)\lambda_{0}}{M^{2}}$. Then $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ implies that
$m=\Omega\left(\frac{1}{(1-\eta)^{2}}\frac{M^{4}B^{4}L(0)}{\lambda_{0}^{4}}\right).$
From Lemma D.4 for the estimation of $L(0)$, i.e.,
$L(0)\lesssim d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right),$
we can deduce that
$m=\Omega\left(\frac{1}{(1-\eta)^{2}}\frac{d^{8}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
(2) When $\sigma$ satisfies Assumption 4.3, we have that
$R\lesssim\frac{\sqrt{(1-\eta)\lambda_{0}}}{d},R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R.$
From Lemma D.4, we can deduce that
$m=\Omega\left(\frac{1}{1-\eta}\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
Proof of Corollary C.1.
Similar as before, when $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{3\lambda_{0}}}{6}$ , we have $\sigma_{min}(\bm{J}(t))\geq\frac{\sqrt{3\lambda_{0}}}{3}$ and then $\lambda_{min}(\bm{H}(t))\geq\frac{\lambda_{0}}{3}$ for $t=0,\cdots,k$.
@@
where the second equality is from the fundamental theorem of calculus and $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)=\bm{w}(t)-s\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(t)$.
Note that $\frac{\partial\bm{u}(\bm{w}(t))}{\partial\bm{w}}=\bm{J}(t)$, thus $\bm{I}_{1}(t)=\eta\bm{u}(t)$. Plugging this into (83) yields that
@@
$\lesssim\frac{\eta\sqrt{L(t)}}{\sqrt{\lambda_{0}}}R^{{}^{\prime\prime}},$
@@
$\leq\left[(1-\eta)^{2}+\frac{C^{2}\eta^{2}(R^{{}^{\prime\prime}})^{2}}{\lambda_{0}}+2(1-\eta)\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\right]\|\bm{u}(t)\|_{2}^{2},$
where $C$ is a universal constant.
Then we can choose $R^{{}^{\prime\prime}}$ such that
$\|\bm{I}_{2}(t)\|_{2}\leq\frac{C\eta\sqrt{L(t)}R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\leq C_{1}\eta\sqrt{L(t)}=C_{1}\eta\sqrt{\bm{u}(t)},$
where $C$ is a universal constant and $C_{1}$ is a constant to be determined.
Thus, we can deduce that
$\|\bm{u}(t+1)\|_{2}^{2}$ $\leq\left[(1-\eta)^{2}+(C_{1}\eta)^{2}+2(1-\eta)C_{1}\eta\right]\|\bm{u}(t)\|_{2}^{2}$
$=\left[(1-\eta)+\eta(\eta C_{1}^{2}+2(1-\eta)C_{1}+\eta-1)\right]\|\bm{u}(t)\|_{2}^{2}$
@@
where in the last inequality is due to that we can choose $C_{1}$ such that $\eta C_{1}^{2}+2(1-\eta)C_{1}+\eta-1\leq 0$.
Note that since $\eta\in(0,1)$, the quadratic equation $\eta x^{2}+2(1-\eta)x+\eta-1=0$ has one negative root and one positive root, denoted as $x_{0}$ and $x_{1}$ respectively. Therefore, the condition $C_{1}\leq x_{1}$ is sufficient to satisfy the requirement. The explicit form of $x_{1}$ can be written as:
$x_{1}=\frac{2(\eta-1)+\sqrt{4(1-\eta)^{2}-4\eta(\eta-1)}}{2\eta}=\frac{\sqrt{1-\eta}}{1+\sqrt{1-\eta}}\geq\frac{\sqrt{1-\eta}}{2}.$
Thus, $C_{1}=\frac{\sqrt{1-\eta}}{2}$ is sufficient to satisfy that $\eta C_{1}^{2}+2(1-\eta)C_{1}+\eta-1\leq 0$.
From this, we can deduce that
$R^{{}^{\prime\prime}}\lesssim C_{1}\sqrt{\lambda_{0}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}.$
Therefore, we can conclude that $\|\bm{u}(t)\|_{2}^{2}\leq(1-\eta)^{t}\|\bm{u}(0)\|_{2}^{2}$ holds for $t=0,\cdots,k$.
10.4. Proof of Corollary 4.9
@@
In the proof of Theorem 4.7, we have proved that Condition 3 holds for all $t\in\mathbb{N}$. Thus, it is sufficient to prove that Condition 3 can lead to the conclusion in Corollary 4.9.
@@
Since $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)$, then for any $r\in[m]$, we have $\|\bm{w}_{r}(s)\|_{2}\leq s\|\bm{w}_{r}(t+1)\|_{2}+(1-s)\|\bm{w}_{r}(t)\|_{2}\leq B$.
When $\sigma(\cdot)$ is smooth, we can deduce that for any $r\in[m]$,
@@
Plugging this into (86), we have
@@
$\lesssim\frac{\sqrt{L(t)}}{\sqrt{\lambda_{0}}}\frac{B^{4}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}$
@@
holds for $t\in\mathbb{N}$, where $C$ is a universal constant.
In the proof above, we only require that $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}=CdR\leq\frac{\sqrt{3\lambda_{0}}}{6}$, leading to the requirement for $m$ that
@@ 11. Auxiliary Lemmas
Lemma 11.1 (Theorem 3.1 in [29]).
If $X_{1},\cdots,X_{n}$ are independent mean zero random variables with $\|X_{i}\|_{\psi_{\alpha}}<\infty$ for all $1\leq i\leq n$ and some $\alpha>0$, then for any vector $a=(a_{1},\cdots,a_{n})\in\mathbb{R}^{n}$, the following holds true:
$P\left(\left|\sum\limits_{i=1}^{n}a_{i}X_{i}\right|\geq 2eC(\alpha)\|b\|_{2}\sqrt{t}+2eL_{n}^{*}(\alpha)t^{1/\alpha}\|b\|_{\beta(\alpha)}\right)\leq 2e^{-t},\ for\ all\ t\geq 0,$
@@
and for $\beta(\alpha)=\infty$ when $\alpha\leq 1$ and $\beta(\alpha)=\alpha/(\alpha-1)$ when $\alpha>1$,
@@
In the following, we will provide some preliminary information about Orlicz norms.
Let $f:[0,\infty)\rightarrow[0,\infty)$ be a non-decreasing function with $f(0)=0$. The $f$-Orlicz norm of a real-valued random variable $X$ is given by
$\|X\|_{f}:=\inf\{C>0:\mathbb{E}\left[f\left(\frac{|X|}{C}\right)\right]\leq 1\}.$
If $\|X\|_{\psi_{\alpha}}<\infty$, we say that $X$ is sub-Weibull of order $\alpha>0$, where
$\psi_{\alpha}(x):=e^{x^{\alpha}}-1.$
Note that when $\alpha\geq 1$, $\|\cdot\|_{\psi_{\alpha}}$ is a norm and when $10<\alpha<$ 1, $\|\cdot\|_{\psi_{\alpha}}$ is a quasi-norm. Moreover, since $(|a|+|b|)^{\alpha}\leq|a|^{\alpha}+|b|^{\alpha}$ holds for any $a,b\in\mathbb{R}$ and $10<\alpha<$ 1, we can deduce that
$\mathbb{E}e^{\frac{|X+Y|^{\alpha}}{|C|^{\alpha}}}\leq\mathbb{E}e^{\frac{|X|^{\alpha}+|Y|^{\alpha}}{|C|^{\alpha}}}=\mathbb{E}e^{\frac{|X|^{\alpha}}{|C|^{\alpha}}}e^{\frac{|Y|^{\alpha}}{|C|^{\alpha}}}\leq\left(\mathbb{E}e^{\frac{2|X|^{\alpha}}{|C|^{\alpha}}}\right)^{1/2}\left(\mathbb{E}e^{\frac{2|Y|^{\alpha}}{|C|^{\alpha}}}\right)^{1/2}.$
This implies that
$\|X+Y\|_{\psi_{\alpha}}\leq 2^{1/\alpha}\max\{\|X\|_{\psi_{\alpha}},\|Y\|_{\psi_{\alpha}}\}\leq 2^{1/\alpha}(\|X\|_{\psi_{\alpha}}+\|Y\|_{\psi_{\alpha}}).$
Furthermore, for $p,q>0$, we have $\||X|\|_{\psi_{p}}=\||X|^{p/q}\|_{\psi_{q}}^{q/p}$. And in the related proofs, we may frequently use the fact that for real-valued random variable $X\sim\mathcal{N}(0,1)$, we have $\|X\|_{\psi_{2}}\leq\sqrt{6}$ and $\|X^{2}\|_{\psi_{1}}=\|X\|_{\psi_{2}}^{2}\leq 6$.
Lemma 11.2.
If $\|X\|_{\psi_{\alpha}},\|Y\|_{\psi_{\beta}}<\infty$ with $\alpha,\beta>0$, then we have $\|XY\|_{\psi_{\gamma}}\leq\|X\|_{\psi_{\alpha}}\|Y\|_{\psi_{\beta}}$, where $\gamma$ satisfies that
$\frac{1}{\gamma}=\frac{1}{\alpha}+\frac{1}{\beta}.$
Proof.
Without loss of generality, we can assume that $\|X\|_{\psi_{\alpha}}=\|Y\|_{\psi_{\beta}}=1$. To prove this, let us use Young’s inequality, which states that
$xy\leq\frac{x^{p}}{p}+\frac{y^{q}}{q},for\ x,y\geq 0,p,q>1.$
Let $p=\alpha/\gamma,q=\beta/\gamma$, then
$\mathbb{E}[\exp(|XY|^{\gamma})]$ $\leq\mathbb{E}\left[\exp\left(\frac{|X|^{\gamma p}}{p}+\frac{|Y|^{\gamma q}}{q}\right)\right]$
$=\mathbb{E}\left[\exp\left(\frac{|X|^{\alpha}}{p}\right)\exp\left(\frac{|Y|^{\beta}}{q}\right)\right]$
$\leq\mathbb{E}\left[\frac{\exp(|X|^{\alpha})}{p}+\frac{\exp(|Y|^{\beta})}{q}\right]$
$\leq\frac{2}{p}+\frac{2}{q}$
$=2,$
where the first and second inequality follow from Young’s inequality. From this, we have that $\|XY\|_{\psi_{\gamma}}\leq\|X\|_{\psi_{\alpha}}\|Y\|_{\psi_{\beta}}$.
Lemma 11.3 (Bernstein inequality, Theorem 3.1.7 in [30]).
@@
Lemma 11.4.
@@
$L(0)=\left\|\begin{pmatrix}\bm{s}(0)\\ \bm{h}(0)\end{pmatrix}\right\|_{2}^{2}=\mathcal{O}\left(d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
@@
$\leq\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right)^{2}+\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}f^{2}(x_{p})$